Instant cold packs, often used to ice athletic injuries on the field, contain ammonium nitrate and water separated by a thin plastic divider. When the divider is broken, the ammonium nitrate dissolves according to the following endothermic reaction: NH4NO

Calculate the change in enthalpy for the reaction in  kJmol−1. (Use 1.0gmL−1 as the density of the solution and 4.184Jg−1∘C−1 as the specific heat capacity.)
Answer #1

q = m c ΔT

where,

m = mass of the substance,

c = specific heat capacity

ΔT = temperature drop

Mass of solution = volume x density = 25.0 mL x 1.0 g/mL = 25.0 g

q = 25.0 g x 4.184 J/goC x (25.7oC − 19.2oC

q = - 679.9 J

q = - 0.6799 kJ

* This answer is negative because we know it is an endothermic reaction (decrease in temp)*

this heat change is being caused by the reaction of 1.50 g of ammonium nitrate.

MW of NH4NO3 = 80 g/mol.

Moles of NH4NO3 = Mass /MW =1.50 g / 80 g/mole = 0.01875 moles


Change in Enthalpy = 0.6799 kJ / 0.01875 mole = 36.26 kJ/mole


Answer to two Significant Figures = 36 kJ/mole


answered by: anonymous